How quickly can you do this? Fill appropriate sign. ( ‘’) 
5 / 7 ☐ 15 / 21

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

5 / 7 = (5 × 3) / (7 × 3)
= 15 / 21
∴5 / 7 = 15 / 21

Thanks